# MSHyper: Multi-Scale Hypergraph Transformer for Long-Range Time Series Forecasting

### v1

@@ Abstract
[…]
1 Introduction
@@
Many time series demonstrate complex and diverse temporal patterns of different scales Wen et al. (2021); Chen et al. (2022, 2023a); Zhang and Yan (2023). For example, due to periodic human activities, traffic occupation and electricity consumption show clear daily and weekly patterns. Considering interactions between these temporal patterns often leads to more accurate forecasting results than analyzing each pattern separately. For example, the morning rush hour on the first workday after a holiday tends to be more congested (interactions between daily and weekly patterns), while the evening rush hour on the last workday before a long holiday starts earlier. Therefore, how to model complex temporal patterns of different scales and their interactions is a fundamental problem in long-range time series forecasting.
@@
To model temporal patterns of different scales and their interactions, traditional methods, e.g., seasonal ARIMA Box and Jenkins (1968) and Prophet Taylor and Letham (2018), use decomposition with heuristic priors to obtain temporal patterns of different scales, but cannot model complex non-linear dependencies of time series. Recently, deep neural networks have demonstrated superiority in capturing non-stationary and non-linear dependencies. Temporal convolutional networks (TCNs) Bai et al. (2018), recurrent neural networks (RNNs) Salinas et al. (2020), and Transformers Zhou et al. (2021); Wu et al. (2021) have been used for time series forecasting. To model temporal patterns of different scales, multi-scale Transformer-based methods Li et al. (2019) attempt to build sub-sequences of different scales from the original input sequence but ignore interactions between temporal patterns of different scales. To address this issue, recent multi-scale Transformer-based methods Liu et al. (2021); Cirstea et al. (2022) introduce special structures (e.g., pyramidal structures) between sub-sequences of different scales. These structures model temporal dependencies within a sub-sequence through intra-scale edges, and model interactions between temporal patterns of different scales through inter-scale edges.
@@
However, these methods only use edges to model pairwise interactions and lack the ability to model high-order interactions (i.e., simultaneous interactions between multiple temporal patterns). In reality, temporal patterns of different scales co-exist and exhibit high-order interactions, e.g., the peak household electricity consumption during summer weekend afternoons (high-order interactions between daily, weekly, and monthly patterns), as well as the high but relatively stable household electricity consumption during winter weekends.
[…]
@@
1) We propose a hypergraph and hyperedge graph construction (H-HGC) module that builds the hypergraph according to the temporal proximity rules, which can model intra-scale, inter-scale, and mixed-scale high-order interactions between temporal patterns. In addition, by treating hyperedges as nodes and building edges based on the sequential relationship and association relationship between nodes, H-HGC also builds the hyperedge graph to enhance hypergraph modeling.
@@
2) We propose a tri-stage message passing (TMP) mechanism that has three message passing phases: Node-hyperedge, hyperedge-hyperedge, and hyperedge-node, which can aggregate pattern information and learn the interaction strength between temporal patterns of different scales.
@@
3) We conduct extensive experiments on five real-world time series datasets, experimental results demonstrate that MSHyper achieves competitive performance, reducing prediction errors by an average of 8.73% and 7.15% over the best baseline in MSE and MAE, respectively.
2 Related Work
@@
Methods for Time Series Forecasting. Time series forecasting methods can be roughly divided into statistical methods and deep neural network-based methods. Statistical methods (e.g., ARIMA Box and Jenkins (1968) and Prophet Taylor and Letham (2018)) follow arbitrary yet simple assumptions, and fail to model complicated temporal dependencies. Recently, deep neural networks show superiority in modeling complicated temporal dependencies. TAMS-RNNs Chen et al. (2021) obtains the periodic temporal dependencies through multi-scale recurrent neural networks (RNNs) with different update frequencies. LSTNet Lai et al. (2018) utilizes recurrent-skip connections in combination with convolutional neural networks (CNNs) to capture long- and short-term temporal dependencies. Transformer-based methods Wen et al. (2022); Chen et al. (2023b) take advantage of the attention mechanism and achieve great access in modeling long-range temporal dependencies. Reformer Kitaev et al. (2020) approximates the attention value through local-sensitive hashing (LSH), Informer Zhou et al. (2021) obtains the dominant query by calculating the KL-divergence to realize the approximate of self-attention, and Autoformer Wu et al. (2021) introduces an auto-correlation mechanism to operate at the level of sub-sequences. However, the above methods struggle to obtain the complex temporal patterns of long-range time series.
@@
Multi-Scale Transformers. Multi-scale or hierarchical Transformers have been proposed in different fields, e.g., natural language processing Ainslie et al. (2020); Guo et al. (2019); Subramanian et al. (2020), computer version Zhang et al. (2021); Dosovitskiy et al. (2021); Wang et al. (2021), and time series forecasting Chen et al. (2022); Zhou et al. (2022); Shabani et al. (2022). Multi-scale ViT Zhang et al. (2021) realizes image recognition by combining multi-scale image feature embeddings and Transformer. Star-Transformer Guo et al. (2019) models intra-scale and inter-scale information interactions by introducing the global node embedding. Due to the limited expression capacity of a single global node, ETC Ainslie et al. (2020) carries out information interactions between the global and local nodes by introducing a set of global node embeddings and setting fixed-length windows. To further extend the ability to model interactions between temporal patterns of different scales, Pyraformer Liu et al. (2021) extends the two-layer structure into multi-scale embeddings, and models interactions between nodes of different scales through a pyramid graph. Crossformer Zhang and Yan (2023) combines a two-stage attention with a hierarchical encoder-decoder architecture to capture cross-time and cross-dimension interactions. However, existing methods only model pairwise interactions between nodes, ignoring high-order interactions between temporal patterns of different scales.
@@
Hypergraph Neural Networks. Hypergraph neural networks (HGNNs) have been proven to be capable of modeling high-order interactions, which have been applied to various fields, e.g., visual object recognition Yan et al. (2020), trajectory prediction Xu et al. (2022), stock selection Sawhney et al. (2021), and citation network classification Bai et al. (2021). HGNN Feng et al. (2019) and HyperGCN Yadati et al. (2019) are the first works to apply graph convolution to hypergraphs, which demonstrate the superiority of hypergraphs over ordinary graph neural networks (GNNs) in modeling high-order interactions. Recent works Yang et al. (2020); Xia et al. (2022) show that HGNNs are promising to model pattern interactions from high-order relations. MBHT Yang et al. (2022) combines hypergraphs with a Transformer framework for the sequential recommendation, which leverages hypergraphs to capture high-order user-item interaction patterns. GroupNet Xu et al. (2022) uses multi-scale hypergraphs for trajectory prediction, which leverages topology inference and representation learning to capture the agent patterns with their high-order interactions.
@@
3 Preliminaries
Hypergraph. A hypergraph is defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is the hyperedge set and $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is the node set. Each hyperedge $e_{m}\in\mathcal{E}$ contains a set of nodes $\{v_{1},v_{2},\ldots,v_{n}\}\subseteq\mathcal{V}$. The hypergraph $\mathcal{G}$ can be denoted as an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where $\mathbf{H}_{nm}=1$ if the $n\text{th}$ node belongs to the $m\text{th}$ hyperedge, else $\mathbf{H}_{nm}=0$. The degree of the $n\text{th}$ node is defined as follows:
@@ 3 Preliminaries
Problem Statement. The task of long-range time series forecasting is to predict the future $H$ steps given the previous $T$ steps of observed values. Specifically, given the input sequence $\mathbb{X}_{1:T}^{\text{I}}=\left\{\bm{x}_{t}\mid\bm{x}_{t}\in\mathbb{R}^{D},t\in[1,T]\right\}$, where $\bm{x}_{t}$ represents the values at time step $t$, and $D$ is the feature dimension. The prediction task can be formulated as:
$\mathbb{X}_{T+1:T+H}^{\text{O}}=\mathcal{F}\left(\mathbb{X}_{1:T}^{\text{I}};\theta\right)\in\mathbb{R}^{H\times D},$ (3)
where $\mathbf{X}_{T+1:T+H}^{\text{O}}$ denotes the forecasting results, $\mathcal{F}$ denotes the mapping function, and $\theta$ denotes the learnable parameters of $\mathcal{F}$.
4 MSHyper
@@
Figure 1: The framework of MSHyper, which consists of three parts: The multi-scale feature extraction (MFE) module, the hypergraph and hyperedge graph construction (H-HGC) module, and the tri-stage message passing (TMP) mechanism.
4.1 Multi-Scale Feature Extraction Module
To get the feature embeddings of different scales, we first map the input sequence into multi-scale embeddings. Concretely, suppose $\mathbf{X}^{s}=\{\bm{x}_{t}^{s}|\bm{x}_{t}^{s}\in\mathbb{R}^{D},t\in[1,h^{s}]\}$ denotes the sub-sequence at scale $s$, where $s=1,...,S$ denotes the scale index, and $S$ is the total number of scales. $h^{s}=\left\lfloor\frac{h^{s-1}}{l^{s-1}}\right\rfloor\text{s.t.}~{}{}s\geq 2$ is the horizon at scale $s$ and $l^{s-1}$ denotes the size of the aggregation window at scale $s-1$. $\mathbf{X}^{1}=\mathbf{X}^{\text{I}}_{1:T}$ is the raw input sequence and the aggregation process can be formulated as:
@@ 4.1 Multi-Scale Feature Extraction Module
where $Aggregation$ is the aggregation function, e.g., 1D convolution or average pooling, and $\theta^{s}$ denotes the learnable parameters of the aggregation function at scale $s$.
4.2 Hypergraph and Hyperedge Graph Construction Module
4.2.1 Hypergraph Construction
Despite existing methods Liu et al. (2021); Wu et al. (2019) being capable of modeling interactions between temporal patterns of different scales through graph structures, we argue that these methods still face two limitations: (1) These methods fail to model high-order interactions between temporal patterns. (2) These methods limit interactions to the neighboring nodes and thus are incapable of capturing interactions between nodes that are far away but still show correlations. To address these problems, we build the hypergraph structures according to the temporal proximity rules. As shown in Figure 2, on the one hand, we build different types of hypergraphs to model intra-scale, inter-scale, and mixed-scale high-order interactions. On the other hand, we build the $k$-hop hypergraph under each type of hypergraph to aggregate information from different ranges of neighbors.
Figure 2: Hypergraph construction. (a), (b), and (c) represent the intra-scale hypergraph, inter-scale hypergraph, and mixed-scale hypergraph, respectively.
Intra-Scale Hypergraph. As shown in Figure 2(a), in order to capture high-order interactions between intra-scale temporal patterns, we construct the intra-scale hypergraph $\mathcal{G}_{\text{ita}}$, which contains two hypergraphs, i.e., the original intra-scale hypergraph $\mathcal{G}_{\text{o, ita}}$ and $k$-hop intra-scale hypergraph $\mathcal{G}_{\text{k, ita}}$. Mathematically, let $\mathcal{G}_{\text{o, ita}}=\{\mathcal{V},\mathcal{E}_{\text{o, ita}}\}$ be the original intra-scale hypergraph, where $\mathcal{E}_{\text{o, ita}}=\{\mathcal{E}_{\text{o, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains original intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e_{i}^{s}\in\mathcal{E}_{\text{o, ita}}^{s}$ is defined as follows:

### v2

@@ Abstract
Demystifying the interactions between temporal patterns of different scales is fundamental to precise long-range time series forecasting. However, previous works lack the ability to model high-order interactions. To promote more comprehensive pattern interaction modeling for long-range time series forecasting, we propose a Multi-Scale Hypergraph Transformer (MSHyper) framework. Specifically, a multi-scale feature extraction (MFE) modlue is introduced to map the input sequence into multi-scale embeddings. Then, a multi-scale hypergraph is designed to provide foundations for modeling high-order pattern interactions and a hyperedge graph is built to enhance hypergraph modeling. In addition, a tri-stage message passing (TMP) mechanism is introduced to aggregate pattern information and learn the interaction strength between temporal patterns of different scales. Extensive experimental results on eight real-world datasets demonstrate that MSHyper achieves the state-of-the-art (SOTA) performance across various settings.
Index Terms:
Time series forecasting, transformer, multi-scale modeling, hypergraph neural networks.
I Introduction
Time series forecasting has demonstrated its wide applications across many fields (e.g., energy consumption planning [57, 21], traffic and economics prediction [20, 27, 50, 62], and disease propagation forecasting [45, 1, 52, 9]. In these real-world applications, how to use a substantial amount of previous time-series data and extend the forecasting horizon into the far future (i.e., long-range time series forecasting) is quite meaningful, as it can help decision makers to make schedule planning and optimize resource allocation.
@@ I Introduction
Many time series demonstrate complex and diverse temporal patterns of different scales [38, 8, 7, 18, 23, 40]. For example, due to periodic human activities, traffic occupation and electricity consumption show clear daily and weekly patterns. Considering the interactions between these temporal patterns often leads to more accurate forecasting results than analyzing each pattern separately. For example, the morning rush hour on the first workday after a holiday tends to be more congested (the interactions between daily and weekly patterns), while the evening rush hour on the last workday before a long holiday starts earlier. Therefore, how to model complex temporal patterns of different scales and their interactions is a fundamental problem in long-range time series forecasting.
@@
To model temporal patterns of different scales and their interactions, traditional methods (e.g., seasonal ARIMA [5] and Prophet [35]), use decomposition with heuristic priors to obtain temporal patterns of different scales, but cannot model complex non-linear dependencies of time series. Recently, deep neural networks have demonstrated superiority in capturing non-stationary and non-linear dependencies. Temporal convolutional networks (TCNs) [3, 29], recurrent neural networks (RNNs) [30, 59, 51], graph neural networks (GNNs) [42, 6, 58, 13], and Transformers [60, 41, 8, 36] have been used for time series forecasting. To model temporal patterns of different scales, multi-scale Transformer-based methods [26] attempt to build sub-sequences of different scales from the original input sequence but ignore the interactions between temporal patterns of different scales. To address this issue, recent multi-scale Transformer-based methods [28, 14] introduce special structures (e.g., pyramidal structures) between sub-sequences of different scales. These structures model temporal dependencies within a sub-sequence through intra-scale edges, and model the interactions between temporal patterns of different scales through inter-scale edges.
@@
However, these methods only use edges to model pairwise interactions and lack the ability to model high-order interactions (i.e., the simultaneous interactions between multiple temporal patterns). In reality, temporal patterns of different scales co-exist and exhibit high-order interactions, e.g., the peak household electricity consumption during summer weekend afternoons (high-order interactions between daily, weekly, and monthly patterns), as well as the high but relatively stable household electricity consumption during winter weekends.
To address the above issue, we propose MSHyper, a Multi-Scale Hypergraph Transformer framework for long-range time series forecasting. MSHyper aggregates the input sequence into sub-sequences of different scales, and models high-order interactions between temporal patterns of different scales by building multi-scale hypergraph structures. To the best of our knowledge, MSHyper is the first work that incorporates hypergraph modeling into long-range time series forecasting. The main contributions are as follows:
@@
  • We propose a hypergraph and hyperedge graph construction (H-HGC) module that builds the hypergraph according to the temporal proximity rules, which can model intra-scale, inter-scale, and mixed-scale high-order interactions between temporal patterns. In addition, by treating hyperedges as nodes and building edges based on the sequential relationship and association relationship between nodes, H-HGC also builds the hyperedge graph to enhance hypergraph modeling.
@@
  • We propose a tri-stage message passing (TMP) mechanism that has three message passing phases: Node-hyperedge, hyperedge-hyperedge, and hyperedge-node, which can aggregate pattern information and learn the interaction strength between temporal patterns of different scales.
@@
  • We conduct extensive experiments on eight real-world time series datasets, and experimental results demonstrate that MSHyper achieves the state-of-the-art (SOTA) performance across various settings.
II Related Work
In this section, we provide a brief review of the related work, including methods for time series forecasting, multi-scale Transformers, and hypergraph neural networks related studies.
II-A Methods for Time Series Forecasting
@@
Time series forecasting methods can be roughly divided into statistical methods and deep neural network-based methods. Statistical methods (e.g., ARIMA [5] and Prophet [35]) follow arbitrary yet simple assumptions, and fail to model complicated temporal dependencies. Recently, deep neural networks show superiority in modeling complicated temporal dependencies [26, 42]. TAMS-RNNs [11] obtains the periodic temporal dependencies through multi-scale recurrent neural networks (RNNs) with different update frequencies. LSTNet [25] utilizes recurrent-skip connections in combination with convolutional neural networks (CNNs) to capture long- and short-term temporal dependencies. Transformer-based methods [39, 12, 24, 60, 41] take advantage of the attention mechanism and achieve great access in modeling long-range temporal dependencies. Reformer [24] approximates the attention value through local-sensitive hashing (LSH), Informer [60] obtains the dominant query by calculating the KL-divergence to realize the approximation of self-attention, and Autoformer [41] introduces an auto-correlation mechanism to operate at the level of sub-sequences. However, the above methods struggle to obtain the complex temporal patterns of long-range time series.
II-B Multi-Scale Transformers
@@
Multi-scale or hierarchical Transformers have been proposed in different fields (e.g., computer vision [55, 16, 37], natural language processing [2, 19, 33], and time series forecasting [8, 61, 32]). Multi-scale ViT [55] realizes image recognition by combining multi-scale image feature embeddings and Transformer. Star-Transformer [19] models intra-scale and inter-scale information interactions by introducing the global node embedding. To address the limitations in the expressiveness of a single global node, ETCformer [2] carries out information interactions between the global and local nodes by introducing a set of global node embeddings and setting fixed-length windows. To further extend the ability to model the interactions between temporal patterns of different scales, Pyraformer [28] extends the two-layer structure into multi-scale embeddings, and models the interactions between nodes of different scales through a pyramid graph. Crossformer [56] combines a two-stage attention with a hierarchical encoder-decoder architecture to capture cross-time and cross-dimension interactions. MSGNet [6] combines an adaptive graph convolution and a temporal multi-head attention mechanism to capture multi-scale inter-series correlations. However, existing methods only model pairwise interactions between nodes, ignoring high-order interactions between temporal patterns of different scales.
II-C Hypergraph Neural Networks
@@
Hypergraph neural networks (HGNNs) have been proven to be capable of modeling high-order interactions, which have been applied to various fields (e.g., visual object recognition [48], sequential recommendation [49, 47], trajectory prediction [44], stock selection [31, 53], and citation network classification [4, 34]). HGNN [17] and HyperGCN [46] are the first works to apply graph convolution to hypergraphs, which demonstrate the superiority of hypergraphs over ordinary graph neural networks (GNNs) in modeling high-order interactions. MBHT [49] combines hypergraphs with a Transformer framework for the sequential recommendation, which leverages hypergraphs to capture high-order user-item interaction patterns. GroupNet [44] uses multi-scale hypergraphs for trajectory prediction, which leverages topology inference and representation learning to capture the agent patterns and their high-order interactions. STHAN-SR [31] leverages hypergraph and temporal Hawkes attention mechanism for stock selection, which leverages hypergraph network architecture to model inter stock relations of varing types and degrees.
@@
III Preliminaries
In this section, we first provide the definition of hypergraph and then formulate the problem.
III-A Hypergraph
A hypergraph is defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is the hyperedge set and $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is the node set. Each hyperedge $e_{m}\in\mathcal{E}$ connects a set of nodes in $\mathcal{V}$. The hypergraph $\mathcal{G}$ can be denoted as an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where $\mathbf{H}_{nm}=1$ if the $n\text{th}$ node belongs to the $m\text{th}$ hyperedge, else $\mathbf{H}_{nm}=0$. The degree of the $n\text{th}$ node is defined as follows:
@@ III-A Hypergraph
III-B Problem Statement
The task of long-range time series forecasting is to predict the future $H$ steps given the previous $T$ steps of observed values. Specifically, given the input sequence $\mathbb{X}_{1:T}^{\text{I}}=\left\{\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t}\in\mathbb{R}^{D},t\in[1,T]\right\}$, where $\boldsymbol{x}_{t}$ represents the values at time step $t$, and $D$ is the feature dimension, the prediction task is defined as follows:
$\widehat{\mathbb{X}}_{T+1:T+H}^{\text{O}}=\mathcal{F}\left(\mathbb{X}_{1:T}^{\text{I}};\theta\right)\in\mathbb{R}^{H\times D},$ (3)
where $\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}$ denotes the forecasting results, $\mathcal{F}$ denotes the mapping function, and $\theta$ denotes the learnable parameters of $\mathcal{F}$.
IV Methodology
In this section, we give the description of the proposed MSHyper. We first give the framework of MSHyper, and then present the detailed descriptions of each module. In addition, we introduce the detailed process of the tri-stage message passing (TMP) mechanism based on the framework of MSHyper.
Figure 1: The framework of MSHyper, which consists of three parts: (a) The HFE module, which maps the input sequence into hierarchical embeddings. (b) The H-HGC module, which provides foundations for modeling high-order interactions between temporal patterns by building the hypergraph and the hyperedge graph. (c) The TMP mechanism, which aggregates pattern information and learns the interaction strength between temporal patterns of different scales.
IV-A Framework
@@
IV-B Multi-Scale Feature Extraction Module
To get the feature embeddings of different scales, we first map the input sequence into multi-scale embeddings. We use $\mathbf{X}^{s}=\{\boldsymbol{x}_{t}^{s}|\boldsymbol{x}_{t}^{s}\in\mathbb{R}^{D},t\in[1,h^{s}]\}$ to represent the sub-sequence at scale $s$, where $s=1,...,S$ denotes the scale index, and $S$ is the total number of scales. $h^{s}=\left\lfloor\frac{h^{s-1}}{l^{s-1}}\right\rfloor\text{s.t.}~{}{}s\geq 2$ is the horizon at scale $s$ and $l^{s-1}$ denotes the size of the aggregation window at scale $s-1$. $\mathbf{X}^{1}=\mathbf{X}^{\text{I}}_{1:T}$ is the raw input sequence and the aggregation process is defined as follows:
@@ IV-B Multi-Scale Feature Extraction Module
where $Aggregation$ is the aggregation function (e.g., 1D convolution or average pooling), and $\theta^{s}$ denotes the learnable parameters of the aggregation function at scale $s$.
IV-C Hypergraph and Hyperedge Graph Construction Module
Long-range time series data contain a lot of noise and are non-stationary [41]. When modeling high-order interactions between temporal patterns without prior knowledge constraints, the model tends to learn spurious interactions and lacks interpretability. To comprehensively model high-order interactions between temporal patterns of different scales, we build the hypergraph and the hyperedge graph separately by the H-HGC module to model intra-scale, inter-scale, and mixed-scale high-order pattern interactions.
IV-C1 Hypergraph Construction
Despite existing methods [28, 43] being capable of modeling the interactions between temporal patterns of different scales through graph structures, we argue that these methods still face two limitations: (1) These methods fail to model high-order interactions between temporal patterns. (2) These methods limit interactions to the neighboring nodes and thus are incapable of capturing the interactions between nodes that are far away but still show correlations. To address these problems, we build the hypergraph structures according to the temporal proximity rules. As shown in Figure 2, on the one hand, we build different types of hypergraphs to model intra-scale, inter-scale, and mixed-scale high-order interactions. On the other hand, we build the $k$-hop hypergraph under each type of hypergraph to aggregate information from different ranges of neighbors.
Figure 2: Hypergraph construction. (a), (b), and (c) represent the intra-scale hypergraph, inter-scale hypergraph, and mixed-scale hypergraph, respectively. In addition to the original connections, we also aggregate information from different ranges of neighbors using $k$-hop connections.
Intra-Scale Hypergraph. As shown in Figure 2(a), in order to capture high-order interactions between intra-scale temporal patterns, we construct the intra-scale hypergraph $\mathcal{G}_{\text{ita}}$, which contains two hypergraphs (i.e., the original intra-scale hypergraph $\mathcal{G}_{\text{o, ita}}$ and $k$-hop intra-scale hypergraph $\mathcal{G}_{\text{k, ita}}$). We use $\mathcal{G}_{\text{o, ita}}=\{\mathcal{V},\mathcal{E}_{\text{o, ita}}\}$ to represent the original intra-scale hypergraph, where $\mathcal{E}_{\text{o, ita}}=\{\mathcal{E}_{\text{o, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains original intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e_{i}^{s}\in\mathcal{E}_{\text{o, ita}}^{s}$ is defined as follows: